OpenStudy (angela210793):

Help please.... x+y=TT Find (1-cosX)/(1+cosy)

OpenStudy (amistre64):

is that a pi symbol perhaps?

OpenStudy (anonymous):

ohhhhhhhhhhhhhhhh

OpenStudy (anonymous):

i was looking for some sort of double angle formula.

OpenStudy (angela210793):

Hmmm...smart :P:P:P:P:P YES IT IS!!! :)

OpenStudy (angela210793):

Oops sorry ^_^

OpenStudy (amistre64):

1-cos(x) ------------ then 1 + cos(pi - x)

OpenStudy (anonymous):

then you get 1, yes?

OpenStudy (amistre64):

you should lol

OpenStudy (angela210793):

OMG!!! That was it? OMG!!! Where's my brain :$:$ :S :(

OpenStudy (anonymous):

\[cos(\pi-x)=cos(\pi)cos(x)+sin(\pi)sin(x)=-cos(x)\]

OpenStudy (angela210793):

Thanks :)

OpenStudy (anonymous):

i wrote that for myself, but it is a basic fact that \[cos(\pi-x)=-cos(x)\]

OpenStudy (angela210793):

yea...I H* trigonometry :(:(

OpenStudy (angela210793):

h8*

OpenStudy (anonymous):

yeah and it only gets worse!